• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name peter

    Log entry time 08:00:43 on August13,2003

    Entry number 110510

    keyword=Owl shift summary

    Shift Leader: Peter Monaghan
    Target Operator: Jing Yuan
    3rd Person: Andrew Puckett

    00:00 Shift begins with MCC trying to restore beam.

    00:45 MCC hope to have beam restored within an hour - latest news from the MCC crew chief.

    01:20 The MCC crew chief has asked if we will accept beam without using the fast feedback mechanism. The corrector for the FFB is being used to compensate for a broken shunt on the Lamberton magnet in the hall A beamline just before the first beam position monitor. If we turn the FFB on, we may break the corrector as it is being worked hard in compensating for the broken shunt. After discussing with our run coordinator, we have agreed to run without the FFB. MCC are now preparing to tune the beam back into the hall.

    01:52 MCC sending 5uA CW beam for harp scan.

    02:05 Starting Harp scans.

    02:51 Harp scan run numbers 2341 - 2359 completed but unfortunately the vertical sigma in both harps is too large (300 - 400 um). MCC are calling an expert for advice to reduce the vertical beam width.

    03:47 Finally finished the Harp scans at run number 2378 (also took a coda run for this scan); results shown below. Changing target to the single carbon longfoil.

    03:53 MCC cycling quadrupole magnets. Given that we only have 4 hours until the energy change, we have decided to change kinematics to 8.26 as we are unsure how long each kinematic will take. We do not believe we have sufficient time to complete two kinematics (8.25 and 8.27) before the beam is removed for the energy change.

    04:20 Kinematic change complete and we are starting with the carbon foil target.

    04:30 Now changing target to the empty reference cell. Y-target for
    the carbon foil is good - 0.2 +/- 0.4mm

    04:53 Empty reference cell run complete; now filling the cell with

    05:03 BMP off in Y as MCC check FFB; restart the N2 reference cell run.

    05:13 N2 reference cell run complete; now changing to the polarised 3He target.

    05:23 Polarised 3He target in place. NMR measurement gives target polarisation = 38.9%. MCC have informed us that the shunt on the Lamberton magnet in front of the first BPM is now working. They are checking the settings and hope to turn the fast feedback on again shortly; beam will be back in about 10 minutes.

    05:33 We have 10uA beam and the fast feedback is now working again.

    06:13 Now we are rotating the target field direction to 270 degrees. MCC have just informed us that they have a vacuum problem in the accelerator which they are investigating. Beam had been off for the previous seven minutes.

    06:27 Target field direction has been rotated to 270 degrees. An update from MCC is that the vacuum problem make take an hour to fix.

    06:36 MCC are bringing hall A to controlled access to allow the vacuum control group to investigate the vacuum problem.

    06:45 MCC have informed us that they believe the vacuum problem may have been caused by beam bleedthrough into hall B, while hall B was in controlled access. They do not believe we will have any beam back before the scheduled energy change this morning.

    08:00 Shift ends with no beam and unfortunately we have been unable to finish kinematic point 8.26.


    Harp Scan # 2378

    Scanner #5: Horizontal = 192um Vertical = 167um
    Scanner #6: Horizontal = 133um Vertical = 185um

    Kinematic: 8.26

    HRS-R/Septum momentum = 0.857
    Septum current = 107.8 A

    04:20 - 04:30 23480/3480 12C foil HWP Right Left

    04:43 - 04:53 23481/3481 Empty Ref. out

    05:00 - 05.02 23482/3482 N2 ref. out BMP off; restart

    05:03 - 05:13 23483/3483 N2 ref. out

    05:34 - 05:53 23484/3484 Pol. 3He out 3M 1.3M

    05.54 - 06:13 23485/3485 Pol. 3He out 2.1M 900k